OpenStudy (anonymous):

(2x^-2+y^-2) divide by (x^-1-y^-1) simplify complex fraction.

OpenStudy (lgbasallote):

\[\Large \frac{2x^{-2} + y^{-2}}{x^{-1} - y^{-1}} \implies \frac{\frac{2}{x^2} + \frac{1}{y^2}}{\frac 1x - \frac 1y} \implies \frac{\frac{2y^2 + x^2}{x^2y^2}}{\frac{y - x}{xy}}\] do you get what i did?

OpenStudy (anonymous):

yes fliped

OpenStudy (lgbasallote):

yup

OpenStudy (anonymous):

gracais

OpenStudy (lgbasallote):

de nada